OpenStudy (anonymous):

Solve by factoring. 5x^2+14x-48=0

OpenStudy (anonymous):

LOL my homework my instructor created. It's a hot mess this week.

OpenStudy (anonymous):

(5x + 24)(x - 2)

OpenStudy (anonymous):

that's the answer :)

OpenStudy (anonymous):

It said incorrecto!

OpenStudy (anonymous):

lol ill play along, wat is the correct answer then :P

OpenStudy (anonymous):

Doesn't give me the answer until I use all three of my chances. It said to do the zero product property/

OpenStudy (anonymous):

u can expand my factorization if u want, it gives u back the original answer and of cause then u can solve for x from here

OpenStudy (anonymous):

x= 2 or -24/5

OpenStudy (anonymous):

I believe after you factor it you have to set each one to 0 so the first one would be 5x+24=0 therefore x= -4.5 second one would be x-2=0 therefore being x=2

OpenStudy (anonymous):

yeah ! I'll give it a try.

OpenStudy (anonymous):

im telling u it is right :P

OpenStudy (anonymous):

ill even argue this with a school teacher :P

OpenStudy (anonymous):

It came up correct. =) I feel like in the morning I am going to have to argue. Good thing it's all online.

OpenStudy (anonymous):

lol xD

OpenStudy (anonymous):

Thank you again, Now off to the next question =)

OpenStudy (anonymous):

ok feel free to mention me and ill help with it =)

OpenStudy (anonymous):

I just posted it

OpenStudy (anonymous):

Find the term that should be added to the expression to form a perfect square trinomial. Write the resulting perfect square trinomial in factored form. x^2+11

OpenStudy (anonymous):

When I do the example it shows it as (11/2)^@= 121/4= x^2+11x+121/4= (x+11/2)^2 but it's saying that's incorrect.